PLEASE PROVIDE ACCURATE STEP-BY-STEP SOLUTIONS.

Question ll The initial velocity of the block in position A is 10 m/s. Knowing that the coefficient of kinetic friction between the block and the plane is 쌔-0.20, determine the time it takes for the block to reach B with zero velocity if θ= 30°
Answer #1

deacceleration = Net force / mass

= mg sin \theta - uk mg cos \theta / m

= g sin \theta - uk g cos \theta

= 9.8 ( sin 30 - 0.2*cos 30)

= 3.2 m/s2

using the equation of linear motion we have

v = u + at

0 = 10 - 3.2*t

t = 3.12 s

so the time taken to reach B is 3.12 s